A pendulum has a period of 2.0s and a amplitude 0.15m.
Calculate,(1).The maximum speed of the pendulum.(2). The maximum acceleration of the Bob
1
Expert's answer
2021-09-16T18:44:04-0400

"\omega=2\pi\nu=\frac{2\pi}T,"

1)

"v=A\omega=0.47~\frac ms,"

2)

"a=A\omega^2=1.48~\frac{m}{s^2}."